# Unifying Two Types of Scaling Laws from the Perspective of Conditional Kolmogorov Complexity

### v1

@@ Abstract
In 2020, OpenAI proposed the first type of Scaling Laws, describing the relationships between model performance and parameters, data, and compute. In 2024, OpenAI proposed the second type of Scaling Laws, describing the relationship between model inference performance and inference computation. In this paper, we analyze LLM training and inference processes from the perspective of lossless compression using conditional Kolmogorov complexity, and unify these two types of Scaling Laws. We find that both types of Scaling Laws improve approximation of conditional Kolmogorov complexity by increasing execution steps $t$. The first type of Scaling Laws increases $t$ by increasing model parameters $y$. The second type of Scaling Laws increases $t$ by increasing the number of output tokens.
@@ 1 Introduction
In October 2022, OpenAI released ChatGPT 3.5 to the public, demonstrating the success of the first type of Scaling Laws proposed in 2020Kaplan et al. (2020). In September 2024, OpenAI released o1-preview and proposed test-time compute Scaling LawsOpenAI (2024). The subsequent series of successful models released by OpenAI have demonstrated the powerful effectiveness of these two types of Scaling Laws.
Hutter (2006) has long advocated that "a model that compresses well generalizes well". Delétang et al. (2023) views large language models as powerful lossless compressors. In this paper, we also analyze large language models from the perspective of lossless compression. However, we differ by adopting the approach from NNCPBellard (2019), viewing the training process of large language models as a form of lossless compression, where the model is trained solely on the data stream to be compressed.
In the realm of theoretical research, the analysis of lossless compression based on Kolmogorov complexity is a prevalent technical approachCover (1985). Employing conditional Kolmogorov complexity as a theoretical tool, we systematically investigates the training process and inference mechanism of large language models.
@@
    We provide a detailed explanation of the relationship between lossless compression and LLM training, reviewing how to train models on and compress the data stream to be compressed. During the compression process, we do not need to transmit the model parameters.
@@
    We model the above compression process using conditional Kolmogorov complexity and prove that training large language models is a computable (total recursive) approximation of the upper bound of joint Kolmogorov complexity. Optimizing joint Kolmogorov complexity naturally has a regularization effect.
@@
    Although Kolmogorov complexity is not computable, we prove that theoretically there exist total recursive functions or transformer neural networks such that $\lim_{t\to\infty}M^{t}(x,y)=C(x\mid y)$
@@
    Through conditional Kolmogorov complexity, we analyze the theoretical limits of large language models in the inference process, showing that there exist infinite problems (strings) that cannot be "solved" by large language models.
@@
    Theoretically, we unify two types of Scaling Laws. Both types of Scaling Laws improve approximation of conditional Kolmogorov complexity by increasing execution steps $t$.
@@ 2 Related Work
Lossless Compression. The brilliant programmer Fabrice Bellard open-sourced NNCPBellard (2019) in 2019, a tool for lossless data compression using neural networks, and upgraded it in 2021 to use autoregressive models (transformers)Bellard (2021) for lossless compression. NNCP trains models on and compresses the data stream to be compressed. Its concept can be traced back to Knoll and de Freitas (2012). In this paper, the authors explained how to achieve lossless compression using machine learning model predictions and arithmetic coding. In fact, lossless compression is equivalent to prediction. Delétang et al. (2023) further advocates viewing large language models as powerful universal lossless compressors, analyzing the first type of scaling laws and the role of tokenization from a lossless compression perspective.
Kolmogorov Complexity. Kolmogorov complexity is closely related to many fundamental principles in machine learning. For example, J. Rissanen’s MDL (Minimum Description Length)Wax and Kailath (1985) can be viewed as a computable approximation of Kolmogorov complexity, while the maximum likelihood principleRissanen (1983) and maximum entropy principleRissanen (1986) can be seen as special cases of the MDL principle. Additionally, this bookLi et al. (2008) further explains the relationships between learning theories, such as PAC learning theory and Kolmogorov complexity. DeMoss et al. (2024) studied the connection between the Grokking phenomenon in neural networks and MDL through Kolmogorov complexity.
@@ 3.1 Turing Machines, Neural Networks & LLMs
The Turing Machine, proposed by British mathematician Alan Turing in 1936, is an abstract computational model designed to explore the computability of problems. As a cornerstone of computer science theory, the Turing Machine has the capability to simulate the execution of any algorithm, providing an important framework for understanding the nature of computation.
From a theoretical perspective, Turing machines are enumerable, meaning all possible Turing machines can be systematically listed:
$T_{1},T_{2},\cdots,$
where each $T_{i}$ represents a unique Turing machine instance. Among these Turing machines, there exists a special class called Universal Turing Machines, denoted as $U$, which can simulate the behavior of any other Turing machine. It’s worth noting that modern computers can be theoretically viewed as physical implementations of Universal Turing Machines, although there are significant differences: modern computers have finite memory, while Universal Turing Machines theoretically have infinite tape.
Neural networks operating on modern electronic computers with finite precision are necessarily equivalent in computational power to some Turing machine $T_{i}$, more precisely, equivalent to a Turing machine that halts in finite steps, i.e., a total recursive function. The process of training neural networks through training data is essentially a process of finding and determining a specific Turing machine $T_{i}$.
Let’s briefly review large language models. Large language models are based on neural network architectures and are trained through "next token prediction" tasks. Given an input sequence (training data) $x_{1:t}=(x_{1},x_{2},\cdots,x_{t})$, the model’s training objective is to predict the next token $x_{t+1}$ in the sequenceBengio et al. (2000). This process can be formalized as the following optimization problem:
$\theta^{*}=\arg\min_{\theta}-\sum_{t=1}^{T-1}\log P(x_{t+1}\mid x_{1:t};\theta)$ (1)
where $P(x_{t+1}\mid x_{1:t};\theta)$ represents the conditional probability of predicting $x_{t+1}$ given the context sequence $x_{1:t}$ under the neural network model with parameters $\theta$, and $T$ represents the sequence length.
From a computational theory perspective, large language models can still be viewed as a special type of Turing machine, specifically one that seeks to minimize 1. The core function of this Turing machine is to calculate the conditional probability distribution of the next token in a predefined vocabulary ${\mathcal{S}}$ based on the given context sequence through complex probability calculations.
@@ 3.2 Dynamic Arithmetic Coding
Dynamic arithmetic codingRissanen (1976)Pasco (1976) is an adaptive data compression algorithm that achieves efficient coding by updating symbol probability distributions in real-time. Here is its mathematical definition:
Let the symbol set be $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{m}\}$, where each symbol $s_{i}$ has probability $P_{t}(s_{i})$ at time $t$, satisfying $\sum_{i=1}^{m}P_{t}(s_{i})=1$.
@@
    Update interval: For input symbol $s_{i}$ at time $t$, update interval to:
    $L_{t}=L_{t-1}+\sum_{j=1}^{i-1}P_{t}(s_{j})\cdot(H_{t-1}-L_{t-1})$
    $H_{t}=L_{t-1}+\sum_{j=1}^{i}P_{t}(s_{j})\cdot(H_{t-1}-L_{t-1})$
@@
    Repeat: Continue processing next symbol until all symbols are processed
@@
    Output: Any number in the final interval as the encoding result. In binary, to achieve shortest encoding, select the decimal with shortest binary representation in the interval.
@@
The decoder reconstructs the original symbol sequence through reverse operations, using the same probability distributions step by step.
It’s important to emphasize that predicted probability distributions $Q_{t}(s_{i})$ can be used for arithmetic coding of symbols. The closeness between $Q$ and the true probability distribution $P$ directly affects coding efficiency - the closer they are, the shorter the average coding length.
Let’s look at a specific example.Consider a symbol set $\mathcal{S}=\{a,b,c,d,e\}$, and we need to encode the message "bab".
$(0.01001)_{2}=(0.28125)_{10}$
@@ 3.3 LLMs & Lossless Compression
In this section, we’ll explore how to achieve lossless data compression using large language models. It’s particularly noteworthy that this method is designed for streaming data, with the core idea being to train the model in real-time using the data stream to be compressed while simultaneously compressing the training data.
@@
Given a large language model $f$, let $\theta_{t}$ represent the model parameters at time $t$. Based on this, the model’s probability distribution prediction for the $(t+1)$-th token can be formally expressed as:
$p_{t+1}=f(x_{t+1}|s_{0},x_{1:t};\theta_{t})$ (2)
where $p_{t+1}$ represents the model’s probability distribution prediction for the $(t+1)$-th token, which is a discrete distribution.
The large language model $f$’s code (such as the Pytorch training code) and various random seeds (such as parameter initializations, etc.) are packaged into a single piece of information. We denote this information as $F$. Now we begin lossless compression of the data $x$. Assume there is an information receiver and an information sender.
@@
    At time 0, execute $f(x_{1}\mid s_{0};\theta_{0})$. Output the prediction probability $p_{1}$ for the first token.
@@
    At time 0, we use arithmetic coding on character $x_{1}$ with $p_{1}$. Note that while $x_{1}$’s true probability distribution $p_{1}^{*}$ is unknown, this does not affect our ability to perform arithmetic coding on $x_{1}$ using the predicted probability distribution $p_{1}$. Note that at this point, arithmetic coding has just begun its first interval division, then selected an interval based on $x_{1}$. We denote this interval as $[l_{1},r_{1}]$.
  3. 3.
    At time 0, perform backpropagation based on the true $x_{1}$ and predicted $p_{1}$ to update $f$’s parameters, obtaining $\theta_{1}$.
  4. 4.
    At time 1, execute $f(x_{2}\mid s_{0},x_{1};\theta_{1})$. Output the prediction probability $p_{2}$ for the second token.
  5. 5.
    At time 1, we use arithmetic coding on character $x_{2}$ with $p_{2}$. Note that at this point, arithmetic coding performs a second interval division based on $[l_{1},r_{1}]$, then selects an interval based on $x_{2}$. We denote this interval as $[l_{2},r_{2}]$. It’s not difficult to see that the entire arithmetic coding process continues, with intervals being continuously subdivided, but the probability $p_{t}$ used for each division changes - this is dynamic arithmetic coding.
  6. 6.
    At time 1, perform backpropagation based on the true $x_{2}$ and predicted $p_{2}$ to update $f$’s parameters, obtaining $\theta_{2}$.
We continuously repeat steps 4 through 6 until time $n-1$. Finally, we will obtain an interval $[l_{n},r_{n}]$ and parameters $\theta_{n}$. We select the shortest binary decimal $z_{n}$ within $[l_{n},r_{n}]$ as the arithmetic coding for the entire information $x$.
@@
  • Arithmetic coding $z_{n}$
  • Code information $F$
  • Required number of decoding iterations $d$
@@
    At time 0, execute $f(x_{1}\mid s_{0};\theta_{0})$. Output the prediction probability $p_{1}$ for the first token.
@@
    At time 0, perform arithmetic decoding based on $p_{1}$ and the received $z_{n}$. The decoding process makes the first interval division based on $p_{1}$. The information receiver will find that $z_{n}$ lies within the interval $[l_{1},r_{1}]$. Thus, the first token $x_{1}$ is decoded.
  3. 3.
    At time 0, perform backpropagation based on the decoded $x_{1}$ and predicted $p_{1}$ to update $f$’s parameters, obtaining $\theta_{1}$. Note that since all random seeds are identical between the information sender and receiver, this $\theta_{1}$ is identical to the previous $\theta_{1}$.
  4. 4.
    At time 1, execute $f(x_{2}|s_{0},x_{1};\theta_{1})$. Output the prediction probability $p_{2}$ for the second token.
  5. 5.
    At time 1, perform arithmetic decoding based on $p_{2}$ and the received $z_{n}$, decoding the second token $x_{2}$. The decoding process makes a second interval division of $[l_{1},r_{1}]$ based on $p_{2}$. The information receiver will find that $z_{n}$ lies within the interval $[l_{2},r_{2}]$. Thus, the second token $x_{2}$ is decoded.
  6. 6.
    At time 1, perform backpropagation based on the decoded $x_{2}$ and predicted $p_{2}$ to update $f$’s parameters, obtaining $\theta_{2}$.
We continuously repeat steps 4 through 6 until time $n-1$ (where $n-1$ is obtained from reading information $d$). Finally, we will have losslessly decoded $z_{n}$ back into the original information $x$. It’s not difficult to see that the entire decoding process did not require prior transmission of the model parameters $\theta_{n}$.
If we use probability distribution $q$ to perform arithmetic coding on information $x$, we can calculate that its average coding length satisfies the following relationship:
$\bar{L}$ $\leq-\sum_{s\in{\mathcal{S}}}p(x=s)\log_{2}q(x=s)+2$ (3)
$=H(p,q)+2$
where $\bar{L}$ denotes the average coding length, $p$ is the true distribution of $x$, and $H(p,q)$ is the entropy between distributions $p$ and $q$. We will use this conclusion to calculate the lossless compression efficiency of large language models.
The challenge in calculating compression rates lies in estimating the coding length of $|z_{n}|$.
For dataset $x$, let ${\hat{p}}_{t}$ represent its true distribution in autoregressive modeling:
${\hat{p}}_{t}=p(x_{t}\mid s_{0},x_{1:t-1})$
@@
$\bar{L}_{z_{n}}$ $\leq\frac{1}{n}(2n+\sum_{t=1}^{n}H({\hat{p}}_{t},p_{t}))$ (4)
$=2+\frac{1}{n}\sum_{t=1}^{n}H({\hat{p}}_{t},p_{t})$
$=2-\frac{1}{n}\sum_{t=1}^{n}\log_{2}{\overline{p}}_{t}$
Note that here ${\overline{p}}_{t}$ is no longer a vector but a value.
The $|F|$ and $|d|$ of the compressed information are basically fixed and negligible compared to $|z_{n}|$. To maximize compression, we need to minimize the average coding length of $z_{n}$. Since directly reducing $z_{n}$ is difficult, and given that we have an upper bound for $z_{n}$, we can instead try to reduce $z_{n}$’s upper bound, namely:
$\min-\frac{1}{n}\sum_{t=1}^{n}\log_{2}{\overline{p}}_{t}$ (5)
Obviously, the optimization objective described in equation 5 exactly matches the training objective of large language models. This indicates that the training process of large language models can essentially be understood as a continuous compression process of the training data.

### v2

@@ Abstract
In 2020, OpenAI proposed the first type of Scaling Laws, describing the relationships between model loss and the scale of parameters, data, and training computation. In 2024, OpenAI proposed the second type of Scaling Laws, describing the relationship between model inference performance and inference computation. In this paper, we analyze LLMs training and inference processes from the perspective of lossless compression using conditional Kolmogorov complexity, and unify these two types of Scaling Laws. We find that both types of Scaling Laws improve approximation of conditional Kolmogorov complexity by increasing execution steps of Turing machine. The first type of Scaling Laws increases execution steps by increasing number of model parameters. The second type of Scaling Laws increases execution steps by increasing the number of intermediate tokens.
@@ 1 Introduction
In October 2022, OpenAI released ChatGPT 3.5 to the public, demonstrating the success of the first type of Scaling Laws proposed in 2020 Kaplan et al. (2020). In September 2024, OpenAI released o1-preview and introduced test-time compute Scaling Laws OpenAI (2024). Subsequently, the release of DeepSeek-R1 Guo et al. (2025) by DeepSeek further validated the effectiveness of test-time compute Scaling Laws. The success of these models highlights the significant impact of both types of Scaling Laws.
Hutter (2006) has long advocated that "a model that compresses well generalizes well". Delétang et al. (2023) views Large Language Models (LLMs) as powerful lossless compressors. In this paper, we also analyze LLMs from the perspective of lossless compression. However, we differ by adopting the approach from NNCP Bellard (2019), viewing the training process of LLMs as a form of lossless compression, where the model is trained solely on the data stream to be compressed.
In the realm of theoretical research, the analysis of lossless compression based on Kolmogorov complexity is a prevalent technical approach Cover (1985). Employing conditional Kolmogorov complexity as a theoretical tool, we systematically investigate the training and inference processes of LLMs.
Figure 1: The left figure achieves more efficient compression of the string $(k,r)$ by increasing the model parameters, while the right figure enhances the compression efficiency of the string $(k,r)$ by introducing more intermediate tokens.
@@
    We provide a detailed explanation of the relationship between lossless compression and LLMs training, reviewing how to train on the stream of data that is to be compressed.
@@
    We model the compression process using conditional Kolmogorov complexity, demonstrating that training LLMs approximates the upper bound of joint Kolmogorov complexity. Optimizing joint Kolmogorov complexity inherently provides regularization.
@@
    Although Kolmogorov complexity is not computable, we prove that theoretically there exist decoder-only transformer such that $\lim_{t\to\infty}M(t,x,y)=C(x\mid y)$.
@@
    Through conditional Kolmogorov complexity, we analyze the theoretical limits of LLMs in the inference process, showing that there exist infinite problems that cannot be "solved" by models.
@@
    We unify two types of Scaling Laws. Both types of Scaling Laws improve approximation of conditional Kolmogorov complexity by increasing execution steps of Turing machine.
@@ 2 Related Work
Lossless Compression. Fabrice Bellard open-sourced NNCP Bellard (2019) in 2019, a tool for lossless data compression using neural networks, and upgraded it in 2021 to use autoregressive models Bellard (2021) for lossless compression. NNCP trains models on the data stream to be compressed. Its idea can be traced back to Knoll and de Freitas (2012). They explained how to achieve lossless compression using machine learning model predictions and arithmetic coding. In fact, lossless compression is equivalent to prediction. Delétang et al. (2023) further advocates viewing LLMs as powerful universal lossless compressors, analyzing the first type of Scaling Laws and the role of tokenization from a lossless compression perspective.
Kolmogorov Complexity. Kolmogorov complexity is closely related to many fundamental principles in machine learning. For example, Minimum Description Length (MDL) proposed by Rissanen (1978) can be viewed as a computable approximation of Kolmogorov complexity, while the Maximum Likelihood Principle Rissanen (1983) and Maximum Entropy Principle Rissanen (1986) can be seen as special cases of the MDL principle. Additionally, Li et al. (2008) further explains the relationships between learning theories, such as PAC learning theory and Kolmogorov complexity. DeMoss et al. (2024) studied the connection between the Grokking phenomenon in neural networks and MDL through Kolmogorov complexity.
@@ 3.1 Turing Machines, Neural Networks & LLMs
The Turing machine Turing (1936), proposed by British mathematician Alan Turing, is an abstract computational model designed to explore the computability of problems. As a cornerstone of computer science theory, the Turing machine has the capability to simulate the execution of any algorithm, providing an important framework for understanding the nature of computation.
From a theoretical perspective, Turing machines are enumerable. Among all Turing machines, there exists a special class called universal Turing machines, denoted as $U$, which can simulate the behavior of any other Turing machine.
Neural networks operating on modern electronic computers with finite precision are equivalent to a Turing machine that halts in finite steps, i.e., a total recursive function. The process of training neural networks through training data is essentially a process of searching for a specific Turing machine.
Next, let’s briefly review LLMs. LLMs are based on neural network architectures and are trained through "next token prediction" tasks. Given an input sequence (training data) $x_{0:t}=[s_{0},x_{1},x_{2},\cdots,x_{t}]$, the model’s training objective is to predict the next token $x_{t+1}$ in the sequence Bengio et al. (2000). Here, $s_{0}$ represents the start token, which is a special token used to indicate the beginning of the sequence. This process can be formalized as the following optimization problem:
$\min_{\theta}-\frac{1}{n}\sum_{t=1}^{n}\log P(x_{t}\mid x_{0:t-1};\theta)$ (1)
where $P(x_{t}\mid x_{0:t-1};\theta)$ represents the conditional probability of predicting $x_{t}$ given the context sequence $x_{0:t-1}$ under the neural network model with parameters $\theta$, and $n$ represents the sequence length.
From a computational theory perspective, LLMs can also be viewed as a special type of Turing machines that minimize Equation 1. The core function of this Turing machine is to calculate the conditional probability distribution of the next token in a predefined vocabulary ${\mathcal{S}}$ based on the given context sequence through complex probability calculations.
@@ 3.2 Dynamic Arithmetic Coding
Dynamic arithmetic coding Rissanen (1976); Pasco (1976) is an adaptive data compression algorithm that achieves efficient coding by updating symbol probability distributions in real-time.
Let the symbol set be $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{m}\}$, where each symbol $s_{i}$ is a sample of a discrete random variable $S$. At time $t$, the probability of symbol $s_{i}$ is denoted by $P_{t}(s_{i}),t=1,2,\cdots,n$, and the probabilities satisfy $\sum_{i=1}^{m}P_{t}(s_{i})=1$.
@@
    Update interval: For input symbol $s_{i_{t}}$ at time $t$, update interval to:
    $l_{t}=l_{t-1}+\sum_{j=1}^{i_{t}-1}P_{t}(s_{j})\cdot(h_{t-1}-l_{t-1})$
    $h_{t}=l_{t-1}+\sum_{j=1}^{i_{t}}P_{t}(s_{j})\cdot(h_{t-1}-l_{t-1})$
@@
    Repeat: Continue processing next symbol in the sequence until all symbols are processed
@@
    Output: Any number in the final interval $[l_{n},h_{n}]$ as the encoding result. In binary, to achieve shortest encoding, select the decimal with shortest binary representation in the interval.
@@
The decoder reconstructs the original symbol sequence through reverse operations, using the same probability distributions step by step. Note that, during the decoding process, apart from the arithmetic encoding information, we also need to know the number of decoding iterations. You can find an example in Appendix A.
It’s important to emphasize that predicted probability distributions $\hat{P}_{t}(s_{i})$ can be used for arithmetic coding of symbol sequence. The closeness between $\hat{P}_{t}$ and the true probability distribution $P_{t}$ directly affects coding efficiency - the closer they are, the shorter the average coding length.
@@ 3.3 LLMs & Lossless Compression
In this section, we’ll explore how to achieve lossless data compression using LLMs.
@@
Given a LLMs $f$, let $\theta_{t}$ represent the model parameters at time $t$. Based on this, the model’s probability distribution prediction for the $(t+1)$-th token can be formally expressed as:
$\hat{p}_{t+1}=f(x_{t+1}|x_{0:t};\theta_{t})$ (2)
where $\hat{p}_{t+1}$ represents the model’s probability distribution prediction for the $(t+1)$-th token, which is a discrete distribution.
The LLMs $f$’s code (such as the Pytorch training code) and various random seeds (such as parameter initializations, etc.) are packaged into a single piece of information. We denote this information as $F$. Now we begin lossless compression of the data $x$. Assume there is an information receiver and an information sender.
@@
    At time 0:
    1. (a)
      Execute $f(x_{1}\mid s_{0};\theta_{0})$. Output the prediction probability $\hat{p}_{1}$ for the first token.
    2. (b)
      Use arithmetic coding on character $x_{1}$ with $\hat{p}_{1}$. Note that while $x_{1}$’s true probability distribution $p_{1}$ is unknown, this does not affect our ability to perform arithmetic coding on $x_{1}$ using the predicted probability distribution $\hat{p}_{1}$. Note that at this point, arithmetic coding has just begun its first interval division, then selected an interval based on $x_{1}$. We denote this interval as $[l_{1},r_{1}]$.
    3. (c)
      Perform backpropagation based on the true $x_{1}$ and predicted $\hat{p}_{1}$ to update $f$’s parameters, obtaining $\theta_{1}$.
@@
    At time 1:
    1. (a)
      Execute $f(x_{2}\mid s_{0},x_{1};\theta_{1})$. Output the prediction probability $\hat{p}_{2}$ for the second token.
    2. (b)
      Use arithmetic coding on character $x_{2}$ with $\hat{p}_{2}$. Note that at this point, arithmetic coding performs a second interval division based on $[l_{1},r_{1}]$, then selects an interval based on $x_{2}$. We denote this interval as $[l_{2},r_{2}]$. It’s easy to see that as the arithmetic coding processes, intervals are continuously subdivided, and the probability $\hat{p}_{t}$ used for each division changes.
    3. (c)
      Perform backpropagation based on the true $x_{2}$ and predicted $\hat{p}_{2}$ to update $f$’s parameters, obtaining $\theta_{2}$.
We continuously repeat steps 2.a through 2.c until $n-1$. Finally, we will obtain an interval $[l_{n},r_{n}]$ and parameters $\theta_{n}$. We select the shortest binary decimal $z_{n}$ within $[l_{n},r_{n}]$ as the arithmetic coding for the entire information $x$.
@@
  • Arithmetic coding $z_{n}$.
  • Program information $F$.
  • Required number of decoding iterations $d$.
@@
    At time 0:
    1. (a)
      Execute $f(x_{1}\mid s_{0};\theta_{0})$. Output the prediction probability $\hat{p}_{1}$ for the first token.
    2. (b)
      Perform arithmetic decoding based on $\hat{p}_{1}$ and the received $z_{n}$. The decoding process makes the first interval division based on $\hat{p}_{1}$. The information receiver will find that $z_{n}$ lies within the interval $[l_{1},r_{1}]$. Thus, the first token $x_{1}$ is decoded.
    3. (c)
      Perform backpropagation based on the decoded $x_{1}$ and predicted $\hat{p}_{1}$ to update $f$’s parameters, obtaining $\theta_{1}$. Note that since all random seeds are identical between the information sender and receiver, this $\theta_{1}$ is identical to the previous $\theta_{1}$.
@@
    At time 1:
    1. (a)
      Execute $f(x_{2}|s_{0},x_{1};\theta_{1})$. Output the prediction probability $\hat{p}_{2}$ for the second token.
    2. (b)
      Perform arithmetic decoding based on $\hat{p}_{2}$ and the received $z_{n}$, decoding the second token $x_{2}$. The decoding process makes a second interval division of $[l_{1},r_{1}]$ based on $\hat{p}_{2}$. The information receiver will find that $z_{n}$ lies within the interval $[l_{2},r_{2}]$. Thus, the second token $x_{2}$ is decoded.
    3. (c)
      Perform backpropagation based on the decoded $x_{2}$ and predicted $\hat{p}_{2}$ to update $f$’s parameters, obtaining $\theta_{2}$.
We continuously repeat steps 2.a through 2.c until $n-1$ (where $n-1$ is obtained from reading information $d$). Finally, we will have losslessly decoded $z_{n}$ back into the original message $x$. In the entire decoding process, the transmission of the model parameters $\theta_{n}$ is not required.
If we use probability distribution $\hat{p}$ to perform arithmetic coding on $S$, we can calculate that its average coding length satisfies the following relationship:
$\bar{L}$ $\leq-\sum_{s\in{\mathcal{S}}}p(s)\log_{2}\hat{p}(s)+2$ (3)
$:=H(p,\hat{p})+2$
where $\bar{L}$ denotes the average coding length, $p$ is the true distribution of $S$, and $H(p,\hat{p})$ is the entropy between distributions $p$ and $\hat{p}$. We will use Equation 3 to calculate the lossless compression efficiency of LLMs.
The challenge of calculating compression rates lies in estimating the coding length $|z_{n}|$, where ||$\cdot$|| ⋅ | denotes the length.
For dataset $x$, let $p_{t}$ represent its true distribution in autoregressive modeling:
$p_{t}=p(x_{t}\mid x_{0:t-1})$
@@
$\bar{L}_{z_{n}}$ $\leq\frac{1}{n}(2n+\sum_{t=1}^{n}H(p_{t},\hat{p}_{t}))$ (4)
$=2+\frac{1}{n}\sum_{t=1}^{n}H(p_{t},\hat{p}_{t})$
$=2-\frac{1}{n}\sum_{t=1}^{n}\log P(x_{t}\mid x_{0:t-1};\theta_{t-1})$
The $|F|$ and $|d|$ are basically fixed and negligible compared to $|z_{n}|$. To maximize compression, we need to minimize the average coding length of $z_{n}$. Since directly reducing $z_{n}$ is difficult, and given that we have an upper bound for $z_{n}$, we can instead try to reduce $z_{n}$’s upper bound, namely:
$\min-\frac{1}{n}\sum_{t=1}^{n}\log P(x_{t}\mid x_{0:t-1};\theta_{t-1})$ (5)
Obviously, the optimization objective described in Equation 5 exactly matches the training objective of LLMs. This indicates that the training process of LLMs can essentially be understood as a continuous compression process of the training data.